OpenStudy (anonymous):

What is the axis of symmetry for the graph of y=2x^2-8x+2 x=?

OpenStudy (anonymous):

is it 3

OpenStudy (anonymous):

x = -2

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

Find algebraically the equation of the axis of symmetry and the coordinates of the vertex of the parabola whose equation is y=-2x^2-8x+3 .. y=-2x^2-8x+3 completing the square y=-2(x^2+4x+4)+3+8 y=-2(x+2)^2+11 Standard form of parabola: y=A(x-h)^2+k, with (h,k) being the (x,y) coordinates of the vertex. A is a multiplier which affects the steepness or narrowness of the parabola. For given parabola, coordinates of the vertex at (-2,11), axis of symmetry, x=-2